# Improved Sample Complexity for Full Coverage in Compact and Continuous Spaces

### v1

@@ Abstract
Verifying uniform conditions over continuous spaces through random sampling is fundamental in machine learning and control theory, yet classical coverage analyses often yield conservative bounds, particularly at small failure probabilities. We study uniform random sampling on the $d$-dimensional unit hypercube and analyze the number of uncovered subcubes after discretization. By applying a concentration inequality to the uncovered-count statistic, we derive a sample complexity bound with a logarithmic dependence on the failure probability ($\delta$), i.e., $M=O(\tilde{C}\ln(\frac{2\tilde{C}}{\delta}))$, which contrasts sharply with the classical linear $1/\delta$ dependence. Under standard Lipschitz and uniformity assumptions, we present a self-contained derivation and compare our result with classical coupon-collector rates. Numerical studies across dimensions, precision levels, and confidence targets indicate that our bound tracks practical coverage requirements more tightly and scales favorably as $\delta\to 0$. Our findings offer a sharper theoretical tool for algorithms that rely on grid-based coverage guarantees, enabling more efficient sampling, especially in high-confidence regimes.
@@ 1 Introduction
However, a significant deficiency of these existing bounds is their linear $1/\delta$ dependence on the failure probability. This means that requiring higher confidence dramatically increases the estimated sample size $M$, often leading to overly conservative estimates. For instance, in the DCDC algorithm context(qu2024deep), when dealing with functions satisfying Lipschitz conditions with constant $\tilde{L}$, the classical approach often overestimates the required sample size, potentially hindering practical implementation in resource-constrained scenarios.
@@
  • We contrast this result with classical coupon-collector-based rates exhibiting linear $1/\delta$ dependence, clarifying when and why the proposed bound yields tighter coverage requirements.
@@ 2 Related Work
Coverage analysis in discrete settings traces back to classical occupancy and coupon collector problems (motwani1996randomized; ferrante2014coupon). In learning theory and Monte Carlo methods, uniform guarantees over discretized spaces are typically established through union bounds or metric entropy arguments (shalev2014understanding; caflisch1998monte; webb2018statistical). In high dimensions, Lipschitz structure and metric entropy guide discretization complexity (valiant1984theory; fournier2011probabilistic). Our work focuses on the uncovered-count statistic and obtains a bound with improved $\ln(1/\delta)$ dependence under uniform sampling; it complements classical expectation-based arguments and aligns with concentration-based perspectives (feller1991introduction). We also connect to verification-style use cases in control-inspired learning where uniform conditions drive correctness (mey2021consistency; qu2024deep).
@@ 3.1 Mathematical Foundations and Core Assumptions
The following core assumptions underpin this study:
@@ 3.2 Main Theorems and Corollaries
Lemma 1:(Expectation of the number of uncovered subcubes)
The expected value of the number of uncovered subcubes $Z$ after $M$ independent samples is:
$E[Z]=\tilde{C}\left(1-\frac{1}{\tilde{C}}\right)^{M}.$ (3)
Proof.
: For any subcube $i$, the probability of being uncovered after $M$ samples is:
$E[Y_{i}]=P(Y_{i}=1)=\left(1-\frac{1}{\tilde{C}}\right)^{M}.$ (4)
By the linearity of expectation, combining equations (2) and (4), we obtain:
@@
Comparing with the classical method, the classic coupon collection model gives $M_{2}=O(\tilde{C}\ln\tilde{C}/\delta)$.
@@ 6 Conclusion and Future Work
This study establishes a tight lower bound of $M=O(\tilde{C}\ln(\frac{2\tilde{C}}{\delta}))$ for complete coverage in uniformly sampled $d$-dimensional unit cubes, representing a fundamental improvement over classical coupon collector bounds. Our analysis leverages spatial discretization and Chebyshev-based probabilistic techniques to achieve logarithmic dependence on failure probability, compared to the classical linear dependence. This improvement is particularly significant for high-confidence scenarios and high-dimensional problems, where classical bounds become overly conservative. The theoretical results are validated through comprehensive numerical experiments, demonstrating practical relevance for algorithm verification and resource allocation. Future work will explore tighter bounds using advanced concentration inequalities and extend the framework to adaptive sampling schemes and broader applications in machine learning and optimization.

### v2

# A Variance-Based Analysis of Sample Complexity for Grid Coverage

@@ Abstract
Verifying uniform conditions over continuous spaces through random sampling is fundamental in machine learning and control theory, yet classical coverage analyses often yield conservative bounds, particularly at small failure probabilities. We study uniform random sampling on the $d$-dimensional unit hypercube and analyze the number of uncovered subcubes after discretization. By applying a concentration inequality to the uncovered-count statistic, we derive a sample complexity bound with a logarithmic dependence on the failure probability ($\delta$), i.e., $M=O(\tilde{C}\ln(\frac{2\tilde{C}}{\delta}))$, which contrasts with the linear $1/\delta$ dependence arising from stopping-time analyses. Under standard Lipschitz and uniformity assumptions, we present a self-contained derivation and compare our result with classical coupon-collector rates. Numerical studies across dimensions, precision levels, and confidence targets indicate that our bound tracks practical coverage requirements more tightly and scales favorably as $\delta\to 0$. Our findings offer a sharper theoretical tool for algorithms that rely on grid-based coverage guarantees, enabling more efficient sampling, especially in high-confidence regimes.
@@ 1 Introduction
However, a notable characteristic of such stopping-time bounds is their linear $1/\delta$ dependence on the failure probability. This means that requiring higher confidence dramatically increases the estimated sample size $M$, often leading to overly conservative estimates. For instance, in the DCDC algorithm context(qu2024deep), when dealing with functions satisfying Lipschitz conditions with constant $\tilde{L}$, the classical approach often overestimates the required sample size, potentially hindering practical implementation in resource-constrained scenarios.
@@
  • We contrast this result with the stopping-time analysis in DCDC (Theorem 5), which exhibits linear $1/\delta$ dependence, and demonstrate that our count-based perspective achieves up to 98.5% sample reduction within the same framework.
@@ 2 Related Work
Coverage analysis in discrete settings traces back to classical occupancy and coupon collector problems (motwani1996randomized; ferrante2014coupon). In learning theory and Monte Carlo methods, uniform guarantees over discretized spaces are typically established through union bounds or metric entropy arguments (shalev2014understanding; caflisch1998monte; webb2018statistical). In high dimensions, Lipschitz structure and metric entropy guide discretization complexity (valiant1984theory; fournier2011probabilistic). Our work focuses on the uncovered-count statistic and obtains a bound with improved $\ln(1/\delta)$ dependence under uniform sampling; it complements classical expectation-based arguments and aligns with concentration-based perspectives (feller1991introduction). We also connect to verification-style use cases in control-inspired learning where uniform conditions drive correctness (mey2021consistency; qu2024deep). We note that DCDC’s Theorem 5 analyzes the stopping time $T$ via $\Pr(T>M)\leq\mathbb{E}[T]/M$, yielding linear $1/\delta$ dependence; our work instead analyzes the uncovered count $Z$ directly, obtaining logarithmic dependence.
@@ 3.1 Mathematical Foundations and Core Assumptions
The following conditions define the problem setting:
@@ 3.2 Main Theorems and Corollaries
By the linearity of expectation, since $E[Y_{i}]=(1-1/\tilde{C})^{M}$ for each subcube $i$, we obtain:
@@
Comparing with DCDC’s stopping-time analysis (Theorem 5 of qu2024deep), which gives $M=O(\tilde{C}\ln\tilde{C}/\delta)$, our count-based bound offers a logarithmic improvement in $\delta$.
@@ 6 Conclusion and Future Work
This study establishes a tight lower bound of $M=O(\tilde{C}\ln(\frac{2\tilde{C}}{\delta}))$ for complete coverage in uniformly sampled $d$-dimensional unit cubes, representing an improvement over stopping-time-based bounds such as DCDC’s Theorem 5. Our analysis leverages spatial discretization and Chebyshev-based probabilistic techniques to achieve logarithmic dependence on failure probability, compared to the classical linear dependence. This improvement is particularly significant for high-confidence scenarios and high-dimensional problems, where classical bounds become overly conservative. The theoretical results are validated through comprehensive numerical experiments, demonstrating practical relevance for algorithm verification and resource allocation. Future work will explore tighter bounds using advanced concentration inequalities and extend the framework to adaptive sampling schemes and broader applications in machine learning and optimization.
@@ Appendix A Appendix: Proof of Theorem 1
The expected number of uncovered subcubes after $M$ samples is $E[Z]=\tilde{C}(1-1/\tilde{C})^{M}$, which follows directly from linearity of expectation.